OpenStudy (anonymous):

A Norman window has the shape of a semicircle atop a rectangle so that the diameter of the semicircle is equal to the width of the rectangle. What is the area of the largest possible Norman window with a perimeter of 30 feet?

OpenStudy (anonymous):

let d = diameter and h = height 30 = (1/2)*d*Pi + d + 2h h = (30 - (1/2)d*Pi-d)/2 = 15 - (1/4)d*Pi - d/2

OpenStudy (anonymous):

area = (1/2)*Pi*(d/2)^2 + d * h

OpenStudy (anonymous):

= (1/2)*Pi*(d/2)^2 + d(15 - (1/4)d*Pi - d/2) = (1/2)*Pi*(d/2)^2 + 15d - (1/4)d^2*Pi-(1/2)d^2

OpenStudy (anonymous):

and at this point i get bored and use mathematica Solve[D[(1/2)*Pi*(d/2)^2 + 15 d - (1/4) d^2*Pi - (1/2) d^2, {d, 1}] == 0, d]

OpenStudy (anonymous):

{{d -> 60/(4 + \[Pi])}}

OpenStudy (anonymous):

(1/2)*Pi*(d/2)^2 + 15 d - (1/4) d^2*Pi - (1/2) d^2 /. d -> 60/(4 + \[Pi]) // N

OpenStudy (anonymous):

63.0112

OpenStudy (anonymous):

thanks!